One number is 12 more that another . The sum of the smaller number and twice the larger number is 39. Find the larger number.
Answer by edjones(8007)

Let the numbers be x and x+12
x+2(x+12)=39
x+2x+24=39
3x+24-24=39-24
3x=15
3x/3=15/3
x=5
x+12=17 Ans.
.
Ed